OpenStudy (anonymous):

if angel 4 = 24 degrees and angel 2 = 101 degrees what is m angle 5?

OpenStudy (anonymous):

lol...angel?

OpenStudy (anonymous):

Angle.. lol

OpenStudy (theviper):

first find angle 3 then u know angle 3,4 so find angle5

OpenStudy (anonymous):

i have no idea what im doing so.. i dont know anything bout angles

OpenStudy (anonymous):

If \(\angle4 + \angle5 = \angle2\), and you put in the values \(24 + \angle5 = 101\), then what is the masure of \(\angle5\)?

OpenStudy (anonymous):

11 66 77 79..

OpenStudy (anonymous):

one of them^

OpenStudy (theviper):

\[\angle2+\angle3=180^o\]\[101+\angle3=180^o\]

OpenStudy (rsadhvika):

lazy lol

OpenStudy (anonymous):

Simplifying further... \(\angle 5 = 101 - 24\)

OpenStudy (anonymous):

77(:

OpenStudy (anonymous):

Good job :)